how do i find the discriminant in the quadratic equation x^2+11x+121=x+96?

Answer by checkley71(8403)

x^2+11x+121=x+96
x^2+11x-x+121-96=0
x^2+10x+25=0
(x+5)^2=0
x+5=0
x=-5 answer.
proof:
-5^2+11*-5+121=-5+96
25-55+121=91
91=91